please help on this: what is the exact value of Tan 5pi/4?
Answer by satyareddy22(84)

Tan 5pi/4
We know that pi=180 degrees
Tan 5pi/4=Tan 5(180)/4=Tan (225)=tan(180+45)=tan 45=1